40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Dec 4 19:34:22 2018, awade, DailyProgress, WOPO, Diabolo 532 nm power drop and retuning 18-12-04_DiaboloInside_annotate.jpegplot20181204_DiaboloSHGRaligned1064nm.pdfplot20181204_DiaboloSHGRErrSig.pdf20181204_RetuneDiaboloSHGAlignment.zip
    Reply  Thu Dec 6 20:04:50 2018, awade, DailyProgress, WOPO, Diabolo 532 nm power drop and retuning 
       Reply  Fri Dec 7 20:31:06 2018, awade, DailyProgress, WOPO, Diabolo: cleaning dust out of SHG unit, self heating issues 18-12-07_20-06-46_4774.jpg18-12-07_20-06-58_4776.jpg18-12-07_20-07-44_4780.jpgIMG_4785.JPGIMG_4766.MOV
          Reply  Thu Dec 13 14:01:13 2018, awade, DailyProgress, WOPO, Diabolo: contacted coherent 
             Reply  Thu Dec 13 16:24:06 2018, awade, DailyProgress, WOPO, Diabolo: baseline numbers for coherent plot20181218_PhaseMatchingCurveFastScanDiaboloSHG.pdf20181213_SHGScanWithLNTempTuning.zip
                Reply  Wed Dec 19 16:48:18 2018, awade, DailyProgress, WOPO, Diabolo: baseline numbers for coherent 
                   Reply  Thu Dec 20 19:20:50 2018, awade, DailyProgress, WOPO, Diabolo: trialing SHG oven set temperature points to see if there is a maxima when locked plot20181220_SHGLockedOutputVsOvenSetpointTemperature.pdf20181220_SHGLockedPowerVsSetPointTemp.zip
                      Reply  Fri Dec 21 15:33:26 2018, Koji, DailyProgress, WOPO, Diabolo: trialing SHG oven set temperature points to see if there is a maxima when locked 
                         Reply  Fri Dec 21 15:51:27 2018, awade, DailyProgress, WOPO, Diabolo: trialing SHG oven set temperature points to see if there is a maxima when locked 
                      Reply  Mon Dec 24 10:08:49 2018, awade, DailyProgress, WOPO, Diabolo: back to power with deliberate pre-misalignment 18-12-21_22-04-30_4812.jpg18-12-21_22-04-54_4813.jpg18-12-21_22-07-55_4814.jpg
                         Reply  Mon Dec 24 12:07:18 2018, Koji, DailyProgress, WOPO, Diabolo: back to power with deliberate pre-misalignment 
Message ID: 2268     Entry time: Tue Dec 4 19:34:22 2018     Reply to this: 2273
Author: awade 
Type: DailyProgress 
Category: WOPO 
Subject: Diabolo 532 nm power drop and retuning 

I restarted the Diabolo laser after it had been shut down for a lab tour.  The using the throlabs S130C head the 1064 nm was 355 mW which is about what we expect (previously it was measured to be 302.3 mW @ 1064 nm. see ATF:2103).  However I was unable to recover the 532 nm full power from the SHG; power from the SHG was down to 5.9 mW (previously ~300 mW).

It seems like mode of the 532 nm is not that clean.  Its likely that the PDH servo is not locking to the optimal mode any more. Because the locked cavity self heats with the high circulating power, often the steady state operation the SHG has a different lock point from the cold start.  Some days I kind of scan it for a bit, lock on the best mode and then come back after a few minutes and relock on the brightest mode.

The last record I made of the Diabolo and its SHG settings was in ATF:2103.  It seems like the doubling crystal temperature has been changed at some stage from 102.87 C to 108.84 C.  We could have done this with Dhruva but it might not have been noted down.  In this original post I noted down values and the SHG was set to its lower bound, this didn't match up with what was in the manual that I had at the time.  Later Koji found the correct manual for this unit (see ATF:2104).  It seems that the manufactures value for peak phase matching was 109.20 C.

Snapshot of settings

Before I mess up the settings by tuning some nobs, here is the present settings for future reference:

Laser diode A temp = 19.69 C
Laser diode B temp = 20.32 C
Injection current  = 2.1 A
Laser crystal temp  = 23.42 C

SHG unit settings were
Double crystal temp = 108.83 C
Offset = 5.09
Gain  = 0.4
Scan amplitude  = 0 (you turn this up to scan SHG cavity)

Fixing alignment of 1064 nm into SHG

I then scanned the SHG cavity to see the transmission peaks and PDH error signal (BNC outputs are on the back of the control unit).  It looked about the same as that reported in ATF:2103.  I then popped the lid on the laser head and tuned the alignment of the two steering mirrors into the SHG to maximize the dominant mode while minimizing the secondary peaks. 

Inside of Diabolo laser head unit.  Steering mirrors were adjusted to optimize alignment.

Trace is displayed below (second attachment).  The lopsidedness of the main peaks are heating effects in the non-linear crystal.  Compare to ATF:2103, ratio of secondary peaks to primary has gone from ~0.25 to <0.05. This should give me a much better power output when relocked.

Still no full power on lock

After replacing lid on laser and attempting to relock the SHG unit I still get the weak mode.  It doesn't look like a HOM but its hard to tell.  

Possible issues are: that we might be getting bad phase matching temperature after locking; self heating is pulling pulling SHG cavity mode to saturate the PZT range; and maybe during the power cycle of the SHG unit, knobs were changed that have messed up the stability of the PDH loop.

I checked the error signal.  Its lopsided, like the transmission peaks, because of the self heating of the Lithium Niobate crystal as we cross resonance. However, zooming in on the oscilloscope seems to show the peak corresponding to the zero crossing point of the error signal. I also played around with the loop gain to find that it is nicely below the point at which it starts ringing.  Its a bit hard to get get an openloop transfer function without getting into the guts of the control box.

SHG PDH signal x10 along with transmission peaks.
Lopsidedness is due to Lithium Niobate self heating, otherwise it looks healthy, if not a bit small.

Bad locking of the SHG is still unresolved.  I will do more troubleshooting today as I can't proceed without 532 nm light.

Attachment 2: plot20181204_DiaboloSHGRaligned1064nm.pdf  123 kB  Uploaded Wed Dec 5 11:19:34 2018  | Hide | Hide all | Show all
plot20181204_DiaboloSHGRaligned1064nm.pdf
Attachment 3: plot20181204_DiaboloSHGRErrSig.pdf  165 kB  Uploaded Wed Dec 5 11:36:34 2018  | Show | Hide all | Show all
Attachment 4: 20181204_RetuneDiaboloSHGAlignment.zip  166 kB  Uploaded Wed Dec 5 12:36:11 2018
ELOG V3.1.3-